# HiTeC: Hierarchical Contrastive Learning on Text-Attributed Hypergraph with Semantic-Aware Augmentation

### v2

@@ A.4 Implementation Details
VilLain (w/o feat.) 74.25 ± 1.90 63.25 ± 2.57 63.18 ± 0.35 65.02 ± 0.23 65.85 ± 0.28 OOM
HyperBERT 51.84 ± 2.03 45.32 ± 1.47 OOM OOM OOM OOM
HiTeC 74.60 ± 1.84 65.91 ± 2.32 71.83 ± 0.21 72.56 ± 0.17 73.91 ± 0.14 67.32 ± 0.09
Table 3: Evaluation results on hyperedge prediction (% mean ± std). The best and second-best results are highlighted in bold and underline, respectively. ‘w/o feat.” means without node features. OOM denotes out-of-memory on a 24GB GPU.
@@
Appendix B Additional Experiment Results
B.1 Hyperedge Prediction Evaluation
B.2 Impact of Text Encoder Backbones
We conduct an ablation study on different text encoder backbones, including DistilBERT (Sanh et al. 2019), BERT-Base (Devlin et al. 2019), RoBERTa-Base (Liu et al. 2019), and DeBERTa-Base (He et al. 2020), as shown in Table 4. Among these, BERT achieves the highest overall performance, followed by DistilBERT. These results suggest that selecting a suitable text encoder for each dataset is important. Nevertheless, our framework is compatible with various backbones. In our main experiments, we adopt BERT as the default text encoder backbone.
B.3 Parameter Sensitivity Analysis of Subgraph Sampling Ratio $r\%$
@@
Appendix C Additional Related Work
C.1 Representation Learning on TAGs
Since hypergraphs can be transformed into graphs via clique expansion, our work is closely related to representation learning on text-attributed graphs (TAGs). Recent TAG methods leverage pre-trained language models to generate numerical features, typically in a cascade (He et al. 2024; Wang et al. 2024) or nested (Yang et al. 2021; Zhao et al. 2023) manner. While most existing TAG methods are supervised, emerging efforts explore self-supervised learning (SSL) to reduce annotation costs. GIANT (Chien et al. 2022a) makes an early attempt by aligning textual and structural features via an extreme multi-label classification objective. More recently, GAugLLM (Fang et al. 2024) proposes an augmentation framework that utilizes large language models (LLMs) to perturb both node features and graph structure, while HASH-CODE (Zhang et al. 2024) formulates a multi-objective contrastive learning strategy tailored for TAGs. However, adapting these methods to hypergraphs via clique expansion incurs substantial structural loss, limiting their effectiveness on TAHGs and motivating the need for dedicated HSSL approaches.

### v3

@@ Appendix D Additional Experiment Results
VilLain (w/o feat.) 74.25 ± 1.90 63.25 ± 2.57 63.18 ± 0.35 65.02 ± 0.23 65.85 ± 0.28 62.92 ± 0.26
HyperBERT 71.84 ± 2.03 64.48 ± 2.10 69.52 ± 0.20 69.34 ± 0.18 70.40 ± 0.15 OOM
HiTeC 75.22 ± 1.87 65.91 ± 2.32 71.83 ± 0.21 72.56 ± 0.17 73.91 ± 0.14 67.32 ± 0.09
Table 4: Evaluation results on hyperedge prediction task(% mean $\pm$ std). The best and second-best results are highlighted in bold and underline, respectively. “w/o feat.” denotes methods without node features. OOM indicates that the method runs out of memory on a single 80GB GPU.
D.1 Hyperedge Prediction Evaluation
D.2 Impact of Text Encoder Backbones
We evaluate different text encoder backbones, including DistilBERT Sanh et al. (2019), BERT-Base Devlin et al. (2019), RoBERTa-Base Liu et al. (2019), and DeBERTa-Base He et al. (2020), as shown in Table 5. Among them, BERT-Base achieves the best overall performance, followed by DistilBERT. The results indicate that backbone selection can noticeably affect performance across datasets. Nevertheless, our framework remains compatible with diverse text encoders. Based on the overall results, we adopt BERT as the default backbone in the main experiments.
@@ D.2 Impact of Text Encoder Backbones
D.3 Parameter Sensitivity Analysis of $s$-walk length $l$
Figure 6: Sensitivity analysis of walk length $l$.
To investigate the effect of walk length on HiTeC, we report the results under different $l\in\{1,2,3,4,5,10\}$ while fixing $s=3$. As shown in Figure 6, model performance generally first improves and then degrades as $l$ increases. This trend suggests that moderate walk lengths help capture richer high-order structural dependencies by exploring broader hypergraph contexts. In contrast, short walks (e.g., $l=1$) provide limited structural context, while excessively long walks (e.g., $l=10$) may introduce noisy or weakly related hyperedges, both leading to suboptimal performance. Overall, $l=4$ achieves a favorable balance between structural exploration and semantic consistency.
D.4 Parameter Sensitivity Analysis of Subgraph Sampling Ratio $r$
@@
Figure 7: Training time and accuracy under different sampling ratios $r$. Bar plots report training time (in seconds), while line plots show the corresponding accuracy. We highlight the relative speedup $\Delta t$ of $r=30$ over $r=50$. $s$ and $l$ are fixed to 3 and 4, respectively. All main results are selected with $r=30$.
D.5 Efficiency Evaluation
We evaluate training time to assess the efficiency of HiTeC. Specifically, we first compare the text encoder pre-training time between HiTeC and HyperBERT, and then compare the hypergraph encoder pre-training time of HiTeC with two representative HCL methods (TriCL and SE-HSSL).
[…]
Method Citeseer Cora History Photo Computers HyperBERT 4h 30m 5h 22m 29h 42m 33h 40m 56h 04m HiTeC 2h 22m 2h 56m 18h 31m 19h 55m 28h 44m
Table 6: Comparison of text encoder pretraining time between HyperBERT and HiTeC.
Hypergraph encoder pretraining time evaluation. We further compare the hypergraph encoder pre-training time of HiTeC with two representative hypergraph-based HCL methods. As shown in Figure 8, training time is plotted on a logarithmic scale for clearer visualization, where HiTeC achieves performance comparable to SE-HSSL while being at least $3\times$ faster than TriCL on large-scale datasets. This efficiency gain mainly stems from two factors: (1) The subgraph-level contrastive objective operates on a small subset of sampled subgraphs (see Appendix D.4); and (2) The preprocessing steps are executed once and reused across multiple training runs. Overall, HiTeC strikes a favorable trade-off between training efficiency and accuracy.
Figure 8: Comparison of hypergraph encoder pretraining time among HiTeC, SE-HSSL, and TriCL. log(·) denotes the natural logarithm.
Appendix E Additional Related Work
E.1 Representation Learning on TAGs
Since hypergraphs can be transformed into graphs via clique expansion, our work is closely related to representation learning on text-attributed graphs (TAGs). Recent TAG methods leverage pre-trained language models to generate numerical features, typically in a cascade He et al. (2024); Wang et al. (2024) or nested Yang et al. (2021); Zhao et al. (2023) manner. While most existing TAG methods are supervised, emerging efforts explore self-supervised learning (SSL) to reduce annotation costs. GIANT Chien et al. (2022a) makes an early attempt by aligning textual and structural features via an extreme multi-label classification objective. More recently, GAugLLM Fang et al. (2024) proposes an augmentation framework that utilizes large language models (LLMs) to perturb both node features and graph structure, while HASH-CODE Zhang et al. (2024) formulates a multi-objective contrastive learning strategy tailored for TAGs. However, hypergraphs fundamentally differ from ordinary graphs in their high-order structural modeling. Converting hypergraphs into ordinary graphs via clique expansion Bron and Kerbosch (1973) inevitably incurs substantial structural information loss Lee and Shin (2023); Li et al. (2024), limiting their effectiveness on TAHGs and motivating the need for dedicated HSSL approaches.